Find all the roots
1. x^2+5x-14=0
2. 3x^2+5x=0
3. 4x^2-13x+3=0
Use the difference of two squares formular to factor
4.y^2-49
5. 4x^2-25

Answer by mananth(16949)

Find all the roots
1. x^2+5x-14=0
x^2+7x-2x-14=0
x(x+7)-2(x+7)=0
(x+7)(x-2)=0
x=-7 OR 2
2. 3x^2+5x=0
x(3x+5)=0
x=0 OR x= -5/3
3. 4x^2-13x+3=0
4x^2-12x-x+3=0
4x(x-3)+1(x-3)=0
(4x+1)(x-3)=0
x=-1/4,x=3
Use the difference of two squares formular to factor
4.y^2-49
(y+7)(y-7)
5. 4x^2-25
(2x+5)(2x-5)